### v1

# Attention-Seeker: Dynamic Self-Attention Scoring for Unsupervised Keyphrase Extraction

Abstract
This paper proposes Attention-Seeker, an unsupervised keyphrase extraction method that leverages self-attention maps from a Large Language Model to estimate the importance of candidate phrases. Our approach identifies specific components – such as layers, heads, and attention vectors – where the model pays significant attention to the key topics of the text. The attention weights provided by these components are then used to score the candidate phrases. Unlike previous models that require manual tuning of parameters (e.g., selection of heads, prompts, hyperparameters), Attention-Seeker dynamically adapts to the input text without any manual adjustments, enhancing its practical applicability. We evaluate Attention-Seeker on four publicly available datasets: Inspec, SemEval2010, SemEval2017, and Krapivin. Our results demonstrate that, even without parameter tuning, Attention-Seeker outperforms most baseline models, achieving state-of-the-art performance on three out of four datasets, particularly excelling in extracting keyphrases from long documents. The source code of Attention-Seeker is available at https://github.com/EruM16/Attention-Seeker.
[…]
Attention-Seeker: Dynamic Self-Attention Scoring for Unsupervised Keyphrase Extraction
Erwin D. López Z.1, Cheng Tang2, Atsushi Shimada2, Graduate School of Information Science and Electrical Engineering, Kyushu University, 1edlopez96s6@gmail.com 2{tang,atsushi}@ait.kyushu-u.ac.jp
1 Introduction
Keyphrase extraction is a critical task in Natural Language Processing (NLP) where the core content of documents is summarized into a set of words or phrases. This process facilitates efficient and accurate information retrieval, which is valuable for several NLP applications, including Retrieval-Augmented Generation (RAG), document categorization, text segmentation, and topic modeling. Keyphrase extraction methods are generally classified into two categories: supervised and unsupervised. Supervised methods achieve high performance by harnessing the power of deep learning techniques, such as neural networks (Song et al., 2021), LSTM Sahrawat et al. (2020), GNN Sun et al. (2019), and transformers Martinc et al. (2022). However, these methods require large amounts of labeled data and are often domain-specific, limiting their practical applicability. In contrast, unsupervised methods rely solely on information extracted from the document itself, making them adaptable across various domains.
Unsupervised keyphrase extraction methods can be divided into five categories based on the type of information they extract. The first two categories include traditional approaches: statistical methods (Spärck Jones, 2004; Campos et al., 2020), which use in-corpus statistical information, and graph-based methods (Mihalcea and Tarau, 2004; Wan and Xiao, 2008; Bougouin et al., 2013; Florescu and Caragea, 2017), which leverage co-occurrence patterns. The remaining three categories consist of more recent works that use Pretrained Language Models (PLMs) to extract semantic information. Embedding-based methods Bennani-Smires et al. (2018); Sun et al. (2020); Ding and Luo (2021); Zhang et al. (2022); Liang et al. (2021) analyze similarities between documents and phrases within the PLM’s embedding space. Prompt-based methods (Kong et al., 2023) use the PLM’s decoder logits to estimate the probability of generating a candidate phrase when a prompt is included in the document (e.g., this book talks about [candidate]). Finally, self-attention-based methods (Kang and Shin, 2023) examine the PLMs’ Self-Attention Map (SAM) to identify which candidate phrases receive the most attention across all document tokens.
As noted by Kong et al. (2023); Kang and Shin (2023), embedding-based approaches struggle to accurately estimate document similarities due to the anisotropic nature of PLM embedding spaces. Prompt and self-attention-based methods circumvent these limitations and achieve state-of-the-art (SOTA) results but introduce new complexities. For example, PromptRank (Kong et al., 2023), a prompt-based method, requires tuning two hyperparameters (alpha and beta) and exploring a wide range of potential prompts. Similarly, SAMRank (Kang and Shin, 2023), a self-attention-based model, requires selecting an optimal SAM from potentially thousands of layer and head combinations. Since documents typically lack labels to guide parameter selection, achieving optimal results without parameterization remains a significant challenge.
In this context, we propose a method that automatically adapts to the characteristics of the input text, thereby eliminating the need for manual parameter tuning. Our approach extends SAMRank (Kang and Shin, 2023), introducing a module that selects the most relevant SAMs to score candidate phrases effectively. Since a direct evaluation of SAMs using labeled keyphrases is impractical, we estimate SAMs’ relevance by defining the characteristics of an optimal SAM. Specifically, Attention-Seeker defines a binary vector that assigns ones to candidate phrases and zeros to all other tokens. It then computes the similarity between this vector and SAMs’ attention vectors (the rows of the self-attention maps), and use the average similarity score to determine the relevance of each SAM. For short documents, Attention-Seeker further evaluates the relevance of individual attention vectors, while for long documents, it considers the relevance of the document segments, refining the attention score estimate for the candidate phrases.
To the best of our knowledge, Attention-Seeker is the first method that automatically adapts its keyphrase extraction process to the specific characteristics of the input text. This novel approach allows Attention-Seeker to achieve SOTA-level performance on four benchmark datasets without requiring any parameter tuning. Our ablation study shows that each of the estimated relevance scores in our method contributes positively to the improvement of a non-parametric SAMRank. Although our proposed relevance scoring mechanism is simple, its effectiveness suggests that further exploration in this direction would be beneficial.
The main contributions of this paper are summarized as follows:
  • We propose Attention-Seeker, a novel unsupervised keyphrase extraction method that automates parameter tuning of self-attention-based models.
  • We propose a simple yet effective approach for selecting the most relevant attention vectors, self-attention maps, and document segments for keyphrase extraction.
  • We demonstrate that non-parametric self-attention-based methods using LLAMA 3-8B achieve high performance on four benchmark datasets, with notable efficacy on long documents.
2 Related work
2.1 Unsupervised Keyphrase Extraction
Traditional unsupervised keyphrase extraction methods can be broadly divided into statistic-based and graph-based approaches. Statistic-based methods estimate word importance using statistical metrics derived from the text. For instance, the TF-IDF method (Spärck Jones, 2004) relies on word frequencies, while YAKE (Campos et al., 2020) incorporates additional factors such as word co-occurrences, positions, and casings. In contrast, graph-based methods represent documents as graphs, with words as nodes, and use graph algorithms to estimate word importance. TextRank (Mihalcea and Tarau, 2004) pioneered this approach by adapting the PageRank algorithm (Page et al., 1998) for keyphrase extraction. Subsequent work has extended this framework: SingleRank (Wan and Xiao, 2008) incorporates information from neighboring documents, TopicRank (Bougouin et al., 2013) clusters nodes within a topic space, and PositionRank (Florescu and Caragea, 2017) integrates a position bias into the PageRank algorithm.
Recent advances use Pretrained Language Models (PLMs) to extract semantic features from documents. These approaches can be categorized into embedding-based, prompt-based, and self-attention-based methods. Embedding-based approaches are the most common, focusing on generating embedding vectors from documents and candidate keyphrases. For example, EmbedRank (Bennani-Smires et al., 2018) uses Doc2Vec (Mikolov et al., 2013) and Sent2Vec (Pagliardini et al., 2018) to derive document and phrase embeddings, using cosine similarity for ranking. SIFRank (Sun et al., 2020) extends this approach by integrating ELMo (Peters et al., 2018) embeddings and the SIF model (Arora et al., 2017) to refine document embeddings. JointGL (Liang et al., 2021) applies a similar strategy using BERT (Devlin et al., 2019) embeddings while incorporating local similarities derived from document graphs. AttentionRank (Ding and Luo, 2021) introduces cross-attention with BERT embeddings to generate document vectors, using self-attention maps (SAMs) for further refinement. Finally, MDERank (Zhang et al., 2022) takes an innovative approach by masking candidate tokens and comparing their similarity to the original document.
Prompt-based methods take a different approach, using PLMs to predict keyphrases directly through prompt conditioning. For example, PromptRank Kong et al. (2023) estimates the likelihood of candidate phrases by using the logits generated by the T5 Raffel et al. (2020) PLM, which is specifically prompted for keyphrase extraction. In contrast, self-attention-based methods, such as SAMRank (Kang and Shin, 2023), focus on using SAMs from PLMs. SAMRank aggregates attention vectors from a specific SAM to measure the attention a candidate phrase receives from other tokens. While SAMRank demonstrates the usefulness of SAMs in keyphrase extraction, it highlights the challenge of effectively selecting the most relevant SAMs without label information.
2.2 Self-Attention Map
Vaswani et al. (2017) introduced the "Scaled Dot-Product Attention" mechanism, as shown in Equation  1. Here, the Query (Q), Key (K), and Value (V) are linear transformations of the embedding representation of the same input sentence. The Softmax function assigns an attention score to the tokens in the Key based on their relevance to a token in the Query, resulting in a matrix known as the Self-Attention Map (SAM). This SAM is used to compute a weighted sum of the Value vectors, which is then transformed into a new embedding representation of the input sentence for further processing by the same attention mechanism in subsequent layers.
$Att\left(Q,K,V\right)=Softmax\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$ (1)
In the transformer model proposed by Vaswani et al. (2017), and in subsequent PLMs, each layer contains multiple heads. As a result, each SAM can focus on different aspects of text processing (multiple heads) at different syntactic levels (multiple layers). Clark et al. (2019) examined the characteristics of different SAMs in the BERT model (Devlin et al., 2019) and identified several types: SAMs that specialize in attending to previous or next tokens, SAMs that focus on separator tokens (e.g., special tokens, periods, and commas), and SAMs that attend to a large number of words (uniform attention). In addition, they identified specific SAMs that preserve syntactic information, suggesting that certain SAMs may also specialize in attending to key topics within the text (Kang and Shin, 2023).
3 Methodology
Figure 1: The core architecture of Attention-Seeker for short documents.
In this section, we introduce our proposed method, Attention-Seeker. It consists of four main steps: (1) Generation of candidate phrases using Part-of-Speech (POS) sequences. (2) Extraction of Self-Attention Maps (SAMs). (3) Estimation of a vector of attention scores from the most relevant SAMs. (4) Scoring the importance of each candidate phrase based on these attention scores. Our method extends SAMRank (Kang and Shin, 2023) by proposing a new approach for estimating the vector of attention scores from all the SAMs (step 3), rather than relying on the manual selection of a single SAM (parameter tuning).
3.1 Candidate Generation
We follow a well-established approach for generating candidate phrases (Kang and Shin, 2023; Ding and Luo, 2021; Bennani-Smires et al., 2018). First, we tokenize and POS tag all words in the document by using the Standford CoreNLP tool [^1] https://stanfordnlp.github.io/CoreNLP/. Next, we extract noun phrases (tagged as “NN”, “NNS”, “NNP”, “NNPS”, and “JJ”) using the NLTK’s RegexpParser [^2] https://github.com/nltk. These phrases are then defined as the document’s keyphrase candidates.
3.2 Extraction of the Self-Attention Maps
We extract the SAMs from each layer and head of the Large Language Model LLAMA 3-8B (AI@Meta, 2024) using the Huggingface library [^3] https://github.com/huggingface/transformers. Similar to SAMRank (Kang and Shin, 2023), we handle the extraction of SAMs differently depending on the document length. For short documents, we input the entire text directly into the LLM. For long documents, we first input the document’s abstract into the LLM, then divide the remaining text into segments with a consistent number of tokens, and process each segment separately.
3.3 Attention Scores Estimation: Short Documents
As shown in Figure 1, the "Attention Score Estimation" module first estimates the relevance of SAMs and their internal attention vectors. These relevance values are then used to compute a weighted average of the attention vectors across all SAMs. Accordingly, the final vector of attention scores $B$ is determined from the most relevant parts of the most relevant SAMs.
To estimate the relevance of each attention vector, we define a hypothesized vector $H$ that characterizes the desired features of a relevant SAM vector. Our initial hypothesis is that SAMs specialized in capturing keyphrases pay attention to tokens belonging to candidate phrases (Equation  2). For instance, SAMs that specialize in attending to token positions or separator tokens are not relevant to our task (Clark et al., 2019).
$H_{i}=\begin{cases}1,&\text{token }i\in\text{candidate}\\ 0,&\text{otherwise}\end{cases}$ (2)
Given a SAM $A^{lh}$, where $l$ is the current layer and $h$ is the current head, the relevance $S^{lh}_{i}$ of each attention vector $A^{lh}_{i}$ is calculated as the dot product between $H$ and $A^{lh}_{i}$. This calculation is formally expressed as the matrix multiplication shown in Equation  3, where $S^{lh}$ is the vector of $S^{lh}_{i}$ values.
$S^{lh}=A^{lh}\cdot H$ (3)
We further set $S^{lh}_{i}$ to 00 for tokens $i$ that do not belong to candidate phrases. We consider their corresponding attention vectors may introduce noise, as they are mapped to the contextualized embeddings of non-candidate tokens in subsequent layers. For example, attention vectors for periods tend to be uniform, helping to create a rich contextual embedding in preparation for a new context.
To determine the relevance $R^{lh}$ of each SAM, we calculate the average relevance score of its individual attention vectors (Equation  4). The final attention vector $B$ is then computed as a weighted average of all attention vectors across all SAMs, with the weights determined by their respective relevance scores (Equation  5).
$R^{lh}=\frac{1}{n}\sum_{i=0}^{n}S^{lh}_{i}$ (4)
$B=\sum_{\forall l,h}\sum_{\forall i}A^{lh}_{i}*S^{lh}_{i}*R^{lh}$ (5)
As shown in Figure  1, the obtained vector $B$ is converted into a new hypothesized vector $H$ by setting to 00 the values $B_{i}$ corresponding to non-candidate tokens. This results in a refined vector of attention scores $B$, now based on the hypothesis that SAMs should focus on tokens belonging to the most important candidate phrases.
3.4 Attention Scores Estimation: Long Documents
Figure 2: The core architecture of Attention-Seeker for long documents.
As described earlier, long documents are segmented into abstract and equally sized segments. Following this segmentation, we estimate attention scores for each segment using the same approach as for short documents (Figure  2). The main difference from our previous method is that relevance scores for attention vectors are not included in the final calculation. This adjustment helps avoid the complexities of applying context-specific relevance scores across multiple, heterogeneous contexts, which could reduce effectiveness without a robust normalization strategy.
Thus, for the segment $s$, the relevance $R^{lh}$ of each of its SAMs is defined as the average of the relevance scores of its corresponding attention vectors (Equation  6). We then apply $l_{1}$-normalization to this vector (Equation  7) to compute a weighted average $SAM^{s}$ of all SAMs from the segment $s$ (Equation  8). This normalization ensures that $SAM^{s}$ preserves the standard properties of a SAM (e.g., the attention values in each row sum to 1). Finally, the attention score vector $B^{s}$ for the segment $s$ is obtained by summing all rows in the average SAM (Equation  9).
$R^{lh}=\frac{1}{n}\sum_{i=0}^{n}A^{lh}_{i}\cdot H$ (6)
$R^{lh}_{n}=\frac{R^{lh}}{\sum_{\forall l,h}R^{lh}}$ (7)
$SAM^{s}=\sum_{\forall l,h}A^{lh}*R^{lh}_{n}$ (8)
$B^{s}=\sum_{\forall i}SAM^{s}_{i}$ (9)
In addition, we define the relevance $T^{s}$ of each segment. To this end, we define a new hypothesized vector $H^{s}$, whose values correspond to the attention scores obtained in the abstract ($B^{0}$) for tokens that belong to the candidate phrases of the abstract, and 0 for all other tokens (Equation  10). Thus, the relevance score $T^{s}$ of the segment $s$ is calculated as the dot product of its attention score vector $B^{s}$ and this newly hypothesized vector (Equation  11).
$H_{i}^{s}=\begin{cases}B^{0}_{j},&\text{token }i=\text{token }j\text{ of }B^{0}\\ 0,&\text{otherwise}\end{cases}$ (10)
$T^{s}=B^{s}\cdot H^{s}$ (11)
3.5 Candidate Final Score Calculation
In short documents, the final score of a candidate phrase is calculated by summing the attention scores $B_{i}$ corresponding to the candidate tokens. Following Kang and Shin (2023), if a candidate phrase consists of a single word, its final score is divided by its frequency. For long documents, we apply the same method to each segment with the vector of attention scores $T^{s}*B^{s}$ ($B^{s}$ multiplied by the relevance of the segment). The final score for each candidate is then obtained by summing its scores across all segments.
Inspec SE2017 SE2010 Krapivin
# Docs 500 493 100 460
Ave. Words 135 194 8154 8545
Ave. Sent. 6 7 380 312
Ave. Keys 9 17 15 6
Unigram (%) 13.5 25.7 20.5 17.8
Bigram (%) 52.7 34.4 53.6 62.2
Trigram (%) 24.9 17.5 18.9 16.4
Table 1: Statistics of datasets
Method Inspec SemEval2017 SemEval2010 Krapivin
F1@5 F1@10 F1@15 F1@5 F1@10 F1@15 F1@5 F1@10 F1@15 F1@5 F1@10 F1@15
Statistic-based Methods
TF-IDF 11.28 13.88 13.83 12.70 16.26 16.73 2.81 3.48 3.91 - - -
YAKE 18.08 19.62 20.11 11.84 18.14 20.55 11.76 14.40 15.19 8.09 9.35 11.05
Graph-based Methods
TextRank 27.04 25.08 36.65 16.43 25.83 30.50 3.80 5.38 7.65 6.04 9.43 9.95
SingleRank 27.79 34.46 36.05 18.23 27.73 31.73 5.90 9.02 10.58 8.12 10.53 10.42
TopicRank 25.38 28.46 29.49 17.10 22.62 24.87 12.12 12.90 13.54 8.94 9.01 8.30
PositionRank 28.12 32.87 33.32 18.23 26.30 30.55 9.84 13.34 14.33 - - -
Embedding-based Methods
EmbedRank d2v 31.51 37.94 37.96 20.21 29.59 33.94 3.02 5.08 7.23 4.05 6.60 7.84
EmbedRank s2v 29.88 37.09 38.40 - - - 5.40 8.91 10.06 8.44 10.47 10.71
[…]
AttentionRank 24.45 32.15 34.49 23.59 34.37 38.21 11.39 15.12 16.66 - - -
MDERank 27.85 34.36 36.40 20.37 31.21 36.63 13.05 18.27 20.35 11.78 12.93 12.58
JointGL 30.82 36.28 36.67 20.49 29.63 34.05 10.78 13.67 14.64 - - -
Prompt-based Methods
[…]
Self-Attention Map-based Methods
[…]
Table 2: Performance of keyphrase extraction of baseline models and Attention-Seeker on four datasets. The bold indicates the best performance and the underline indicates the two best performances.
4 Experiments and Results
4.1 Datasets and Evaluation Metrics
We evaluate Attention-Seeker using four benchmark datasets: Inspec (Hulth, 2003), SemEval2010 (Kim et al., 2010), SemEval2017 Augenstein et al. (2017), and Krapivin (Krapivin et al., 2009). These datasets consist of scientific papers and vary in length. Specifically, Inspec and SemEval2017 contain paper abstracts (short documents), while SemEval2010 and Krapivin contain full papers (long documents). Table  1 presents more detailed statistical information about each dataset.
The performance of keyphrase extraction is evaluated with the F1 score at the top 5 (F1@5), 10 (F1@10), and 15 (F1@15) keyphrases predicted by the model (candidates with the highest final scores). The final set of predicted keyphrases is determined after applying the NLTK’s PorterStemmer for stemming and removing duplicate candidates.
4.2 Baselines
Our baselines include statistics-based methods: TF-IDF (Spärck Jones, 2004), YAKE (Campos et al., 2020); graph-based methods: TextRank (Mihalcea and Tarau, 2004), SingleRank (Wan and Xiao, 2008), TopicRank (Bougouin et al., 2013), PositionRank (Florescu and Caragea, 2017); embedding-based methods: EmbedRank (Bennani-Smires et al., 2018), SIFRank (Sun et al., 2020), AttentionRank (Ding and Luo, 2021), MDERank (Zhang et al., 2022), JointGL (Liang et al., 2021); prompt-based methods: PromptRank (Kong et al., 2023), and self-attention-based models: SAMRank (Kang and Shin, 2023). To ensure a fair comparison, we implemented SAMRank using the average of all SAMs (a non-parametric method) extracted with the LLAMA 3-8B model. The optimized results of SAMRank are detailed in Appendix  A. Additionally, we omitted the proportional score, as it led to lower performance in this setting (Appendix  B). Consequently, the reported results of SAMRank represent the most important baseline for our model, as they correspond to the same method when all relevance scores are assigned equal values.
Method Inspec SemEval2017 SemEval2010 Krapivin
F1@5 F1@10 F1@15 F1@5 F1@10 F1@15 F1@5 F1@10 F1@15 F1@5 F1@10 F1@15
Base 34.25 38.18 38.11 24.74 33.51 37.01 16.80 20.26 20.94 16.38 16.44 14.78
$B+S^{lh}$ 34.98 39.30 38.99 25.02 34.38 38.14 15.77 20.37 21.43 16.70 16.55 15.07
$B+R^{lh}$ 34.44 38.54 38.24 24.85 33.79 37.11 16.79 20.60 21.26 16.99 16.48 14.91
$B+f$ 34.80 39.54 39.08 25.00 34.30 37.63 15.90 19.80 21.00 16.59 16.67 15.24
$B+S^{lh}+R^{lh}$ 35.07 39.56 39.05 25.25 34.57 38.17 16.63 20.41 21.36 16.92 16.58 15.25
$B+S^{lh}+R^{lh}+f$ 35.23 40.22 39.48 25.27 34.78 38.33 15.83 20.40 21.52 16.64 16.72 15.45
Attention-Seeker 35.49 40.14 39.22 25.40 34.53 38.50 16.74 20.26 21.73 17.14 16.89 15.27
Table 3: Performance of Attention-Seeker using different relevance scores in short documents.
4.3 Results
Table  2 summarizes the performance of Attention-Seeker compared to baseline models across four benchmark datasets. The results show that Attention-Seeker achieves state-of-the-art (SOTA) performance on the Inspec, SemEval2010, and Krapivin datasets, while securing the second-best performance on the SemEval2017 dataset. The non-parametric SAMRank also performs well on all datasets, highlighting the effectiveness of Self-Attention Maps (SAMs) for unsupervised keyphrase extraction. In particular, the performance of self-attention-based methods on long document datasets (SemEval2017 and Krapivin) suggests a promising direction for extracting keyphrases from long documents using SAMs from LLMs. In this context, the consistent outperformance of Attention-Seeker over SAMRank demonstrates the advantages of considering the relevance of different SAMs and document segments when estimating the attention scores.
Although PromptRank remains the best-performing method on the SemEval2017 dataset, it requires hyperparameter tuning for each specific benchmark (Kong et al., 2023), which limits its suitability for applications without labeled data. Furthermore, Kong et al. (2023); Song et al. (2024) demonstrate that the choice of prompts significantly affects its performance, requiring an extensive search to identify the optimal prompt. In contrast, Attention-Seeker achieves the second-best performance on this dataset without requiring parameter tuning. As Kang and Shin (2023) note, different input texts activate different SAMs for keyphrase identification. Attention-Seeker adapts to these variations by analyzing the SAMs triggered by the input, positioning it as a robust and adaptable solution for unsupervised keyphrase extraction.
5 Ablation study
5.1 Short documents
Attention-Seeker integrates several modules designed to improve its overall performance. Specifically, it employs a relevance score for each SAM’s attention vector ($S^{lh}$), a relevance score for each SAM ($R^{lh}$), and a filtering step $f$ applied to non-candidate tokens of $S^{lh}$ (Section  3.3). To evaluate their contributions, we applied these modules separately. We also tested configurations using both scores $S^{lh}$ and $R^{lh}$, with and without the filtering step. Table  3 presents the performance of Attention-Seeker in all these cases. In the "Base" configuration, all attention vectors and SAMs are assigned equal relevance, and no filtering is applied. The following configurations progressively incorporate the relevance scores ($S^{lh}$ and $R^{lh}$) and the filtering step ($f$) into the base method ($B$). The final version of Attention-Seeker corresponds to "$B+S^{lh}+R^{lh}+f$" applied twice (Figure  1). Performance on the SemEval2010 and Krapivin datasets is evaluated using only the abstracts of the papers (short documents).
Table  3 shows that each module contributes positively to the overall performance of our method, with the $S^{lh}$ score providing the most significant improvements. While our proposed relevance scores are supported by the observations of Clark et al. (2019), the benefit of filtering attention vectors from certain tokens is less straightforward. Clark et al. (2019) observed that in SAMs specialized in attending to verbs of direct objects, attention vectors of non-nouns often attend to the [SEP] token. They suggest that in such cases, this token might function as a "non-op", implying that the remaining attention scores would be randomly distributed (the model ignores this information). Our results in the "$B+f$" configuration suggest this hypothesis is correct, as the filtering step would be removing these noisy scores. However, results from SemEval2010 indicate that our filtering approach may also cause slight performance degradation, likely due to the simplicity of the current filtering mechanism. Therefore, future research should investigate how SAM’s attention vectors differ based on different query tokens and determine which tokens produce more effective attention distributions for representing contextual information.
Method SemEval2010 Krapivin
F1@5 F1@10 F1@15 F1@5 F1@10 F1@15
Base 17.86 20.99 22.07 17.38 16.78 15.15
$B_{as}$ 17.35 21.04 21.05 17.05 16.50 14.83
$B_{as}+R^{lh}$ 17.34 21.27 21.83 17.44 16.79 15.17
$B_{as}+T^{s}_{b}$ 18.20 21.52 22.72 18.58 17.39 15.88
$B_{as}+T^{s}$ 19.05 22.75 23.69 20.49 17.94 16.01
Attention-Seeker 19.00 23.07 23.81 20.79 18.25 16.22
Table 4: Performance of Attention-Seeker using different relevance scores in long documents.
5.2 Long documents
Attention-Seeker segments the document into abstract and equal-length segments of the remaining content. Since this approach slightly differs from the base method (Kang and Shin, 2023; the entire document is segmented into equal parts), we evaluated its impact on the final performance. In addition, we evaluated the contributions of the relevance score for each SAM ($R^{lh}$) and the relevance score for each document segment $T^{s}$. We considered two versions of this second relevance score: One uses our proposed refined vector $H^{s}$ ($T^{s}$), and the other uses a simple binary vector $H^{0}_{b}$ that assigns one to abstract candidates and zero otherwise ($T^{s}_{b}$). Table  4 summarizes the results. In the "Base" and "$B_{as}$" configurations, all SAMs and segments are treated as equally relevant ("$B_{as}$" considers the abstract as the first segment). The "$B_{as}$" configuration is then extended with relevance scores ($R^{lh}$, $T^{s}_{b}$, and $T^{s}$), and the final Attention-Seeker corresponds to $B_{as}+R^{lh}+T^{s}$.
Table  4 shows that our two proposed relevance scores contribute positively to the performance of Attention-Seeker, with the relevance of segments providing the most significant improvements. In addition, refining the hypothesized vector $H^{s}$ further improves the performance of the segment scores ($T^{s}>T^{s}_{b}$). Despite these advances, our segmentation approach introduces some performance degradation compared to the base model, likely due to the difference in length between the abstract and other segments. While the overall improvement achieved by our method outweighs the initial degradation, further research may consider exploring new methods for estimating segment relevance without compromising the performance of the base model.
Finally, the contribution of SAMs’ relevance scores $R^{lh}$ in Table  4 is consistent with the results in Table  3, suggesting that our approach for short documents is equally effective when applied to segments of long documents. This result suggests that the inclusion of the relevance score $S^{lh}$ has the potential to further improve the performance of Attention-Seeker for long documents. Further research should explore methods for normalizing attention scores across contexts that vary in syntax, number of nouns, topics, and other factors.
5.3 Relevance of SAMs
Attention-Seeker’s score $R^{lh}$ identifies the SAMs most relevant for keyphrase extraction, eliminating the need for labeled data to select an optimal SAM. Our analysis of the $R^{lh}$ scores across samples from four datasets (Inspec, SemEval2017, SemEval2010, and Krapivin) reveals that SAMs from intermediate layers (9 to 15 out of 32), along with a few SAMs from the first and last layer contribute significantly to the keyphrase extraction task (see Appendix  C).
6 Conclusion
We proposed Attention-Seeker, a self-attention-based method for unsupervised keyphrase extraction that eliminates the need for manual parameter tuning. Our method assumes that certain Self-Attention Maps (SAMs) are specialized to focus on the most important phrases within a document. Accordingly, Attention-Seeker reframes keyphrase extraction as the task of identifying where the most crucial information is encapsulated in SAMs. For long documents, we extended this method by first identifying the most relevant segments from which to extract attention scores.
Attention-Seeker outperformed most baselines on four benchmark datasets, demonstrating the effectiveness of our approach. Since Attention-Seeker is the first method to estimate the relevance from the internal scores of a Large Language Model (LLM), its scoring mechanism is very simple and has potential for improvement, as suggested in our study. Future research should explore more sophisticated methods for relevance estimation, which could lead to further performance improvements. In addition, optimizing these relevance estimates may provide deeper insights into the internal processes of LLMs, potentially contributing to the design of more efficient pretrained language models and their application in unsupervised tasks.
7 Limitations
Attention-Seeker estimates the relevance of attention vectors by defining a set of desired characteristics and measuring how closely each attention vector matches them. However, the current implementation of this approach present two major limitations. First, it uses the dot product between the desired vector and SAMs’ attention vectors. While effective, this approach could be improved by applying normalization techniques such as $l_{2}$-normalization (cosine similarity), softmax normalization (cross-attention), or softmax with temperature scaling. Second, the hypothesized vectors ($H$), which serve as reference vectors, are initially defined as binary vectors that attend to candidate phrases. This simplistic formulation may introduce noise into the attention distribution of the candidates. Although we propose a refinement of this vector ($H$) using the attention scores extracted from the document ($B$), the initial estimation of $B$ still relies on the original binary vector. To address these limitations, future work could explore more sophisticated hypothesized vectors and more effective methods for measuring their alignment with SAM’s attention vectors.
Another important limitation is that the long document version of Attention-Seeker requires the first segment to be the abstract of the document. This condition limits the application of our proposed method to documents with an abstract. In addition, our ablation study showed that this approach negatively affects the overall performance of our method. Accordingly, we believe that further research should explore alternative methods for defining the hypothesized vector $H^{s}$ without relying on information extracted from the abstract. Possible solutions could include extracting the most important tokens from all segments under the assumption of equal relevance, and defining the hypothesized vector $H^{s}$ based on these tokens.
Appendix A Optimal Performance of SAMRank on LLAMA 3-8B
We adapted the original SAMRank method (Kang and Shin, 2023) from the official repository https://github.com/kangnlp/SAMRank, using the LLAMA 3-8B model instead of GPT-2. Due to the structural similarities between GPT-2 and LLAMA 3-8B, no major modifications were required. Table  5 shows the results obtained by the optimal SAMs on four benchmark datasets. We evaluated the performance of the method using both Global ($G$) and Final ($F$) scores.
Dataset F1@5 F1@10 F1@15 SAM
SAMRank Final Score (S)
Inspec 33.19 39.04 39.50 L:19 H:15
SemEval2017 24.50 34.05 37.65 L:21 H:11
SemEval2010 16.30 18.04 19.74 L:12 H:2
Krapivin 18.55 17.25 15.63 L:8 H:8
SAMRank Global Score (G)
Inspec 35.60 40.47 40.21 L:9 H:7
SemEval2017 25.64 36.41 39.80 L:20 H:28
SemEval2010 17.10 19.32 20.40 L:12 H:2
Krapivin 19.18 17.60 15.93 L:12 H:2
Table 5: Performance of keyphrase extraction of SAMRank (Global and Final scores) on four datasets and their corresponding heads.
The performance of the final score in SAMRank LLAMA 3-8B is consistent with the results reported by Kang and Shin (2023) for SAMRank using GPT-2, LLAMA 2-7B and LLAMA 2-13B. However, in our implementation, the independent Global Score provides the best results, suggesting that the Proportional Score may be less effective in this context. While the ablation study conducted by Kang and Shin (2023) demonstrated the benefits of the Proportional Score for BERT and GPT-2 implementations, its effectiveness for larger LLMs such as LLAMA remains unclear. We believe that further research is needed to determine the impact of the Proportional Score on LLM implementations of SAMRank.
The independent implementation of the Global Score in SAMRank outperforms Attention-Seeker in short documents, suggesting room for improvement in our method. As discussed in the main paper, these improvements could include more effective methods for defining the desired characteristics of attention vectors and better measuring their alignment with these characteristics. While we might expect results similar to or better than those reported in Table  5 (Global Score), we must also consider the possibility that non-relevant SAMs might achieve high performance by chance due to the use of non-ideal labels and a large set of 1024 trials.
Finally, Attention-Seeker outperforms the Global Score of SAMRank in long documents, highlighting the importance of our proposed relevance score for document segments $T^{s}$. While this score proves crucial for extracting keyphrases from long documents, the results also respond to the limitation of SAMRank in selecting a single SAM for all document segments. When different segments have their own optimal SAM, this constraint leads to suboptimal results. In addition, the reduced likelihood of selecting a non-relevant SAM contributes to a more stable selection with lower performances. This stability is reflected in the results of the SemEval2010 and Krapivin datasets in Table  5, where in 3 out of 4 cases, the selected SAM is consistently from layer 12, head 2.
Appendix B Results of Non-Parametric SAMRank
The non-parametric SAMRank introduced in this study implements the method proposed by SAMRank (Kang and Shin, 2023). However, instead of selecting a single SAM from the heads of the LLM (as in the original parameterized method), it averages all SAMs. Given the set of SAMs $A=\{A^{lh},\forall l,h\}$, the averaged SAM $A^{av}$ is calculated as shown in Equation  12 ($p$=$q$=32 for LLAMA 3-8B).
$A^{av}=\frac{1}{p*q}\sum_{l=1}^{p}\sum_{h=1}^{q}A^{lh}$ (12)
The global score is calculated as shown in Equation  13 ($n$ is the number of tokens of the document or segment).
$G=\sum_{i=1}^{n}\sum_{h=1}^{m}A^{av}_{ij}$ (13)
The proportional score is calculated as shown in Equation  14. Here, $A^{x}$ is equal to the matrix $A^{av}$, where its elements $A^{av}_{i0}$ are set to zero.
$P=\sum_{j=1}^{n}\frac{A^{x}_{ij}}{\sum_{i=1}^{n}A^{x}_{ij}}$ (14)
Equation  14 is a simplification of the equations proposed by Kang and Shin (2023), as their redistribution of the global score $G$ is cancelled out by the column normalization of the matrix. The only observable effect of this redistribution is that all elements $A^{av}_{i0}$ become zero, since the global score $G$ for the first token is zero.
The final score $S$ is determined by the sum of the global score $G$ and the proportional score $P$, as shown in Equation  15. The score for each candidate phrase is derived from the attention scores of this vector, as detailed by Kang and Shin (2023).
$S=G+P$ (15)
Dataset F1@5 F1@10 F1@15
SAMRank Global Score (G)
Inspec 34.25 38.18 38.11
SemEval2017 24.74 33.51 37.01
SemEval2010 17.86 20.99 22.07
Krapivin 17.38 16.78 15.15
SAMRank Proportional Score (P)
Inspec 27.21 33.81 35.20
SemEval2017 20.49 25.90 33.01
SemEval2010 15.75 19.07 19.34
Krapivin 14.84 14.45 13.27
SAMRank Final Score (S)
Inspec 30.23 36.62 38.08
SemEval2017 21.95 31.83 35.39
SemEval2010 16.36 20.13 20.87
Krapivin 16.06 15.46 14.29
Table 6: Performance of keyphrase extraction of the three scores of a non-parametric SAMRank on four datasets.
We implemented the non-parametric SAMRank method using the model LLAMA 3-8B, evaluating their three different scores to ensure that the non-parametric version benefits from these scores similarly to the original SAMRank. Table  6 shows the performance achieved by the non-parametric SAMRank in the three cases.
We observe a similar result than in the optimized implementation of SAMRank. The Proportional Score degrades the improvement of the overall performance, causing the implementation with the Global Score to achieve the best results. Accordingly, the performance reported in Table  2 is collected from the implementation of SAMRank that only use the Global Score.
Appendix C Most Relevant SAMs identified by Attention-Seeker
Figure 3: Relevance of LLAMA 3-8B layers for keyphrase extraction estimated by Attention-Seeker over four datasets. (The lighter the color, the higher the relevance rank of the corresponding layer)
Attention-Seeker estimates the relevance of SAMs, their internal attention vectors, and their corresponding segments. While analyzing the relevance of attention vectors and document segments requires a deep study of different cases (influence of both the input text and the selected SAM), the relevance of SAMs is more closely related to the model characteristics. Specifically, the layers of LLMs represent different levels of text processing, with higher layers capturing more complex token relationships. To investigate the impact of different layers on keyphrase extraction, we extracted the relevance scores $R^{lh}$ estimated by Attention-Seeker across all samples of the four datasets used in our study (Inspec, SemEval2017, SemEval2010, and Krapivin).
Figure  3 shows the distribution of $R^{lh}$ scores for each layer, with the scores normalized for consistent comparison. For better visualization, the top 5% of outliers are filtered out from each distribution, distributions are color-coded according to their relevance rank, and a red line is added at $Relevance=3.125$ to indicate a uniform distribution where all layers are equally relevant. The analysis shows that the heads between layers 8 and 18, especially layers 9, 12, 14, and 15, received the highest relevance scores. This result suggests that heads located in intermediate layers of an LLM tend to specialize in processing tasks related to noun-noun relationships and selecting the most relevant nouns/phrases.
The first and last layers show high variability in relevance scores across heads, with some receiving remarkably high scores. This suggests that attention to keyphrases is a fundamental task: At low levels, it helps to generate well-contextualized embeddings, while at high levels, it contributes to building a complete representation of the document context. At both levels, however, this process is complemented by numerous other tasks performed by different heads.
Figure 4: Relevance of LLAMA 3-8B heads for keyphrase extraction estimated by Attention-Seeker in one sample of the Inspec dataset.
Figure 5: Attention scores in Layers 3 and 14 in one sample of the Inspec dataset.
Figure  4 supports the previous observation by illustrating the relevance scores $R^{lh}$ assigned to each head for a sample from the Inspec dataset. The score distribution across layers aligns with the general pattern in Figure  3, where most heads in the intermediate layers contribute significantly to keyphrase extraction, while only a few heads in the first and last layers exhibit high relevance. For the same Inspec sample, we compared the attention scores of candidate phrase tokens in layer 3 (lowest overall relevance score) and layer 14 (highest overall relevance score). Figure  5 shows the difference between these two layers: Layer 14 assigns higher attention scores to all candidates, with a clearer contrast between phrases, while layer 3 tends to focus on non-candidate tokens, resulting in lower, more uniformly distributed attention scores.